OpenStudy (anonymous):

There are 2 brown chocolates and 3 white chocolates in a bag. You select a chocolate and roll a dice ( 1-6). What is the probability of getting a brown chocolate and roll a 6 on the dice?

OpenStudy (anonymous):

these are independent events abd thr probabiliyies are multiplied P(brown choc) = 2/5 P(6) = 1/6 so reqd. prob. = 2/5 * 1/6 = 2/30 = 1/15

OpenStudy (anonymous):

these are independent events abd thr probabiliyies are multiplied P(brown choc) = 2/5 P(6) = 1/6 so reqd. prob. = 2/5 * 1/6 = 2/30 = 1/15